# Order Parameter Discovery for Quantum Many-Body Systems

### v1

@@ Appendix B Solution of the order parameter discovery problem
Appendix C Additional results
C.1 Discontinuities of fidelity and its susceptibility
We produce a counterexample which proves that the assumption on the non-degeneracy of the ground state (stated in Section III) is not sufficient to guarantee the continuity of fidelity when dealing with RDM. For example, consider a parametric state |$\ket{\psi(\theta)}\in\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, with $\mathcal{H}_{A}=\mathbb{C}^{2}$ and $\mathcal{H}_{B}=\mathbb{C}^{n}$, defined by the Schmidt decomposition [43]
[…]
with {|$\{\ket{\varphi_{k}}\}$ and {|$\{\ket{b_{k}}\}$ denoting some bases for $\mathbb{C}^{2}$ and $\mathbb{C}^{n}$, respectively. We obtain the RDM by tracing out the subsystem $B$, so
$\rho(\theta)=$ $\operatorname{Tr}_{B}(\ket{\psi(\theta)}\bra{\psi(\theta)})$
$=$ $\cos^{2}(\theta)\ket{\varphi_{1}}\bra{\varphi_{1}}+\sin^{2}(\theta)\ket{\varphi_{2}}\bra{\varphi_{2}}.$ | . (40)
It is clear that the parameter $\theta\in\mathbb{R}$ determines a change of rank of $\rho(\theta)$. In the case of a RDM acting on $\mathbb{C}^{2}$, the Uhlmann fidelity matches the formulation of the Superfidelity [44] which is defined as
$F_{S}(\rho,\sigma)\coloneqq$ $\operatorname{Tr}(\rho\sigma)+\sqrt{\left(1-\operatorname{Tr}\left(\rho^{2}\right)\right)\left(1-\operatorname{Tr}\left(\sigma^{2}\right)\right)}.$ (41)
Then considering $F_{S}(\rho(\theta),\rho(\theta+h))$, the rightmost term for the fidelity in (41) becomes
$c\sqrt{\left(1-\operatorname{Tr}\left(\rho^{2}(\theta+h)\right)\right)}=$ $c\left|\sin(\theta+h)\cos(\theta+h)\right|,$ (42)
where $c$ is a constant depending on $\theta$. We see that the resulting function of $h$ is discontinuous for $h+\theta\in\{k\pi/2|k\in\mathbb{Z}\}$. In other words, a tiny perturbation could cause a change of rank of $\rho(\theta)$ and a loss of continuity (and so smoothness).
In practice, we do not know the locations in the parameters space where such phenomena could happen. However, the procedure of sampling and smoothing described in Appendix A can be interpreted as an implicit mollification. Consequently, the plot of the mollified vector field (for example Figure 4) should present sharp changes in direction, sources or sink in the proximity of such regions.
Appendix D Experimental Methods
@@
To calculate the ground states for the reduced fidelity susceptibility, we use the density matrix renormalization algorithm (DMRG) [45]. We use two different software packages called Tensor Network Python TeNPy [35] and Quantum Simulation with MPS Tensor qs-mps [46].
@@
For the ANNNI model we span the region $(\kappa,h)\in\mathcal{R}=[0.01,1.5]\times[0.01,1.5]$ while for the cluster $(K,h)\in\mathcal{R}=[0.5,1.5]\times[0.5,1.5]$. We take $n=64$ points for each axis resulting in a $64\times 64$ lattice of parameters, with a maximum bond dimension $\chi=64$. Both Hamiltonians in the top-left corner of their respective regions $\mathcal{R}$ are dominated by the $\sigma^{z}$ term and thus the states will be a perturbation of the all-up state $|\ket{\uparrow\uparrow\cdots\uparrow}$. To ensure the convergence of the DMRG calculations, we use the following strategy: We begin with a region where the states are relatively easy to prepare and where the DMRG converges reliably. From there, we extend the calculation to include nearest-neighbor lattice points, using the previously obtained state as the initial state. We continue this approach, gradually moving towards regions where states are more challenging to prepare. In doing so, we gain both accuracy and speed. We note the most time consuming part of the calculation is solving the local eigenvalue problem through the method eigsh from the software package SciPy [47]. This procedure is effective up until the point of a phase transition. At that point, if the computation time exceeds a specified threshold (which can be set either arbitrarily or adaptively), we shelve the calculation and proceed to the next lattice point.
Appendix E Discovered Order Parameter of the ANNNI Model
@@
$+0.217\cdot ZZ$
Appendix F The ANNNI model
@@
$h_{I}\approx\frac{1-\kappa}{\kappa}\left(1-\sqrt{\frac{1-3\kappa+4\kappa^{2}}{1-\kappa}}\right)$ (43)
@@
$h_{KT}(\kappa)\approx 1.05\sqrt{\left(\kappa-\frac{1}{2}\right)(\kappa-0.1)}$ (44)
@@
as highlighted in [52] and corresponds to the green line in Figure 3. We also note the presence of the Lifshitz point, marked by a red dot in Figure 3. This point represents a critical point, known as the Lifschitz point (LP), on the phase diagram of certain magnetic systems, where a line of second-order phase transitions meets a line of incommensurate modulated phases [55]. Lastly, we observe that the ANNNI model features a disorder line, known as the Peschel-Every (PE) line [56], which lies within the paramagnetic phase and provides a reference point for comprehending the properties of this particular phase.
@@
Finally, we note this model also reproduces important features observed experimentally in systems that can be described by discrete models with effectively short-range competing interactions [15]. These experimental findings include Lifshitz points [57, 58], adsorbates, ferroelectrics, magnetic systems, and alloys. Conversely, the so-called floating phase emerging in the model is appealing to experimental researchers to explore. This critical incommensurate phase has been observed very recently by using Rydberg-atom ladder arrays [59].

### v3

@@ Appendix B Solution of the order parameter discovery problem
Appendix C Special solution of the order parameter discovery problem
In this section, we obtain the solution for the special case of the order parameter discovery established by the problem in (IV). The outcome is the close form solution put forward in (17).
Fixed the density matrices $\rho_{+},\rho_{-}$ of order $m$, we define the linear operator $\Xi:\mathcal{S}_{m}\to\mathcal{S}_{m}$ on the space of Hermitian matrices, given by the rule
$\Xi(K)\coloneqq$ $-\frac{\left\langle K,\rho_{+}\right\rangle}{\left\langle\rho_{+},\rho_{+}\right\rangle}\rho_{+}+\frac{\left\langle K,\rho_{-}\right\rangle}{\left\langle\rho_{-},\rho_{-}\right\rangle}\rho_{-},$ (45)
with $K\in\mathcal{S}_{m}$. Let $C\in\mathcal{S}_{m}$ be any non-zero Hermitian, then the operator
$K\mapsto\frac{\left\langle K,C\right\rangle}{\left\langle C,C\right\rangle}C$ (46)
is a projection onto the 1-dimensional subspace spanned by $C$. Consequently, the operator in (45) can be related to the Gram-Schmidt process (GS) [31]. We obtain the quadratic form for the operator $\Xi$, so for $K\in\mathcal{S}_{m}$
$\langle K,\Xi(K)\rangle=$ $-\frac{\operatorname{Tr}\left(\rho_{+}K\right)^{2}}{\operatorname{Tr}(\rho_{+}^{2})}+\frac{\operatorname{Tr}\left(\rho_{-}K\right)^{2}}{\operatorname{Tr}(\rho_{-}^{2})}\in\mathbb{R}.$ (47)
The latter is the objective of the problem in (IV), so the optimization can be rewritten as
$\underset{M\in\mathcal{S}_{m}}{\min}$ $\langle M,\Xi(M)\rangle,$ (48)
s.t. $\|M\|_{F}^{2}\leq 1.$
We obtain the solution of the latter problem by means of the singular value decomposition (SVD) of the operator $\Xi$, which is established by the following result.
Theorem C.1.
Let $\Xi:\mathcal{S}_{n}\to\mathcal{S}_{n}$ be the linear operator defined in (45). Assume that for the density matrices $\rho_{+},\rho_{-}$, defining $\Xi$, it holds that $\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\rangle\neq 1$. Then,
$\Xi(K)=$ $\|V_{1}\|_{F}\left\langle K,\widehat{V_{1}}\right\rangle U_{1}+\|U_{2}\|_{F}\left\langle K,V_{2}\right\rangle\widehat{U_{2}}$ (49)
$=$ $\sqrt{1-\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\rangle^{2}}\cdot\left(\left\langle K,\widehat{V_{1}}\right\rangle U_{1}+\left\langle K,V_{2}\right\rangle\widehat{U_{2}}\right)$ (50)
is the reduced-SVD of $\Xi$, with non-zero singular values
$s_{1}=s_{2}=\sqrt{1-\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\rangle^{2}},$ (51)
and unnormalized right $V_{i}$ and left $U_{i}$ singular vectors
$V_{1}=$ $\widehat{\rho_{+}}-\left\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\right\rangle\widehat{\rho_{-}};\quad U_{1}=-\widehat{\rho_{+}},$ (52a)
$V_{2}=$ $\widehat{\rho_{-}};\quad U_{2}=\widehat{\rho_{-}}-\left\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\right\rangle\widehat{\rho_{+}}.$ (52b)
Proof.
We start by rewriting the operator $\Xi$ defined in (45) in the equivalent form
$\Xi(K)=$ $-\left\langle K,\widehat{\rho_{+}}\right\rangle\widehat{\rho_{+}}+\left\langle K,\widehat{\rho_{-}}\right\rangle\widehat{\rho_{-}}.$ (53)
Consider the operator
$P(K)=$ $\left\langle K,\widehat{\rho_{-}}\right\rangle\,\left\langle\widehat{\rho_{-}},\widehat{\rho_{+}}\right\rangle\,\widehat{\rho_{+}},$ (54)
then adding and subtracting the latter to $\Xi$ we obtain
$\Xi(K)=$ $\left(-\left\langle K,\widehat{\rho_{+}}\right\rangle+\left\langle K,\widehat{\rho_{-}}\right\rangle\,\left\langle\widehat{\rho_{-}},\widehat{\rho_{+}}\right\rangle\right)\cdot\widehat{\rho_{+}}$ (55a)
$+\left\langle K,\widehat{\rho_{-}}\right\rangle\cdot\left(\widehat{\rho_{-}}-\left\langle\widehat{\rho_{-}},\widehat{\rho_{+}}\right\rangle\,\widehat{\rho_{+}}\right)$
$=$ $\left\langle K,\widehat{\rho_{+}}-\left\langle\widehat{\rho_{-}},\widehat{\rho_{+}}\right\rangle\widehat{\rho_{-}}\right\rangle\cdot\left(-\widehat{\rho_{+}}\right)+\left\langle K,V_{2}\right\rangle U_{2}$
$=$ $\left\langle K,V_{1}\right\rangle U_{1}+\left\langle K,V_{2}\right\rangle U_{2},$
with $V_{i}$ and $U_{i}$ defined in (52a). In general, we have that $\langle\rho_{+},\rho_{-}\rangle\neq 0$, so $K$ can have a non-zero projection on both terms of (53). However, in (55a) we can immediately verify that $\langle V_{1},V_{2}\rangle=0$ and $\langle U_{1},U_{2}\rangle=0$. In addition, we verify that
$\langle V_{1},V_{1}\rangle=$ $\langle U_{2},U_{2}\rangle=1-\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\rangle^{2}\in(0,1],$ (56)
thus we can rewrite (55a) as stated in (50). Hence, the expression in (50) is the reduced-SVD of $\Xi$, as stated in the claim. ∎
As the final step, we identify the close formulation for the solution of the problem in (48), using the SVD of the operator $\Xi$. We assume the conditions of Theorem C.1. Then, by substituting the right singular vectors into the quadratic form in (47) we obtain
$\left\langle\widehat{V_{1}},\Xi(\widehat{V_{1}})\right\rangle=$ $1-\frac{\operatorname{Tr}\left(\rho_{+}\rho_{-}\right)^{2}}{\operatorname{Tr}(\rho_{+}^{2})}>0,$ (57a)
$\left\langle\widehat{V_{2}},\Xi(\widehat{V_{2}})\right\rangle=$ $-\left(1-\langle\widehat{\rho_{+}},\widehat{\rho_{-}}\rangle^{2}\right)<0.$ (57b)
Hence $M=\widehat{V_{2}}$ (expanded in (17)) is the observable that minimizes the objective in (47), subject to the norm of the Hermitian $M$ being 1. In Appendix D we showcase the exact formula applied to the Ising model.
Appendix D Exact example for the order parameter of the Ising model
We consider the quantum Ising model in a transverse magnetic field and the representative density matrices $\rho_{+},\rho_{-}$ for its phases: if the interaction term dominates or the magnetic field strength dominates, respectively. We denote with $\sigma^{x},\sigma^{y},\sigma^{z}$ the corresponding Pauli matrices. Let $\rho_{+}=\frac{\mathbb{I}+\sigma^{z}}{2}$ and $\rho_{-}=\frac{\mathbb{I}+\sigma^{x}}{2}$. Both $\rho_{+}$ and $\rho_{-}$ are norm 1 orthogonal projections, and $\langle\rho_{+},\rho_{-}\rangle=\frac{1}{2}$. We apply the formula for the optimal observable in (17) to get
$M^{\star}=$ $\frac{\mathbb{I}+2\sigma^{z}-\sigma^{x}}{2\sqrt{3}}=\frac{\sigma^{z}}{\sqrt{3}}+\frac{\mathbb{I}-\sigma^{x}}{2\sqrt{3}}$ (58)
which is a norm 1 Hermitian. Clearly, for the disordered phase $\operatorname{Tr}(\rho_{-}M^{\star})=0$, whereas $\operatorname{Tr}(\rho_{+}M^{\star})=\frac{\sqrt{3}}{2}$ for the ordered phase. Moreover, we note that $M^{\star}$ is a linear combination of the orthogonal components $\sigma^{z}$ and $I-\sigma^{x}$. The $\sigma^{z}$ operator is the local order parameter or scaling operator, while $I-\sigma^{x}$ measures the magnetization in the $x$-direction.
Now, let $M=\sigma^{z}/\|\sigma^{z}\|_{F}=\sigma^{z}/\sqrt{2}$ (norm 1 as for $M^{\star}$), that is the observable $M$ is the magnetization in the $z$-direction for present model. Then, $\operatorname{Tr}(\rho_{-}M)=0$ and $\operatorname{Tr}(\rho_{+}M)=\frac{1}{\sqrt{2}}$, hence we confirm that
$\operatorname{Tr}(\rho_{+}M)\leq\operatorname{Tr}(\rho_{+}M^{\star}).$ (59)
In other words, we have shown that the observable $M^{\star}$ presents an increased gap between the phases, when compared to the known magnetization $M$ or it maximizes the value in the order phase.
Appendix E Experimental Methods
@@
To calculate the ground states for the reduced fidelity susceptibility, we use the density matrix renormalization algorithm (DMRG) [62]. We use two different software packages called Tensor Network Python TeNPy [54] and Quantum Simulation with MPS Tensor qs-mps [63]. In the case of the Rydberg model instead, we use ITensor [64].
@@
For the ANNNI model we span the region $(\kappa,h)\in\mathcal{R}=[0.01,1.5]\times[0.01,1.5]$ while for the cluster $(K,h)\in\mathcal{R}=[0.5,1.5]\times[0.5,1.5]$. We take $n=64$ points for each axis resulting in a $64\times 64$ lattice of parameters, with a maximum bond dimension $\chi=64$. Both Hamiltonians in the top-left corner of their respective regions $\mathcal{R}$ are dominated by the $\sigma^{z}$ term and thus the states will be a perturbation of the all-up state $|\ket{\uparrow\uparrow\cdots\uparrow}$. To ensure the convergence of the DMRG calculations, we use the following strategy: We begin with a region where the states are relatively easy to prepare and where the DMRG converges reliably. From there, we extend the calculation to include nearest-neighbor lattice points, using the previously obtained state as the initial state. We continue this approach, gradually moving towards regions where states are more challenging to prepare. In doing so, we gain both accuracy and speed. We note the most time-consuming part of the calculation is solving the local eigenvalue problem through the method eigsh from the software package SciPy [65]. This procedure is effective up until the point of a phase transition. At that point, if the computation time exceeds a specified threshold (which can be set either arbitrarily or adaptively), we shelve the calculation and proceed to the next lattice point.
Appendix F Discovered Order Parameter of the ANNNI Model
@@
$+0.217\cdot ZZ.$
Appendix G The ANNNI model
@@
$h_{I}\approx\frac{1-\kappa}{\kappa}\left(1-\sqrt{\frac{1-3\kappa+4\kappa^{2}}{1-\kappa}}\right).$ (60)
@@
$h_{KT}(\kappa)\approx 1.05\sqrt{\left(\kappa-\frac{1}{2}\right)(\kappa-0.1)}.$ (61)
@@
as highlighted in [70] and corresponds to the green line in Figure 3. We also note the presence of the Lifshitz point, marked by a red dot in Figure 3. This point represents a critical point, known as the Lifschitz point (LP), on the phase diagram of certain magnetic systems, where a line of second-order phase transitions meets a line of incommensurate modulated phases [73]. Finally, we note the presence of a disordered line in the ANNNI model, referred to as the Peschel-Every (PE) line [74] within the paramagnetic phase. This line serves as a reference point for understanding the characteristics of this specific phase.
@@
We note this model also reproduces important features observed experimentally in systems that can be described by discrete models with effectively short-range competing interactions [17]. These experimental findings include Lifshitz points [75, 76], adsorbates, ferroelectrics, magnetic systems, and alloys. Conversely, the so-called floating phase emerging in the model is appealing to experimental researchers to explore. This critical incommensurate phase has been observed very recently by using Rydberg-atom ladder arrays [77].
Appendix H Vector field patterns
$x_{1}$$x_{2}$$\mathbf{c}(t)$$\mathbf{c}(t_{0})$$\mathbf{c}^{\prime}(t_{0})$$\nabla_{\mathbf{x}}g(\mathbf{c}(t_{0}))$
Figure 18: Curve $\mathbf{c}(t)$ whose image belongs to a level set of a function $g:\mathcal{X}\subseteq\mathbb{R}^{2}\to\mathbb{R}$, that is $g(\mathbf{c}(t))$ is constant for $t$ in some open interval.
This appendix aims to highlight the patterns that emerge in the plot of the vector fields. Specifically, we focus on the ”source”-like patterns observed at the phase transitions and the ”sink”-like patterns that become evident. We demonstrate the conditions under which these patterns appear in the plot of the vector field $P(\boldsymbol{\lambda})$ (defined in (12)).
Let $g:\mathcal{X}\subseteq\mathbb{R}^{n}\to\mathbb{R}$ be a differentiable function and define the level set [78]
$L(k)\coloneqq$$\left\{\mathbf{x}\middle|\mathbf{x}\in\mathcal{X},g(\mathbf{x})=k\right\}$ (63)
for some $k\in\mathbb{R}$, as the subset of the domain of $g$ whose image under $g$ is the constant $k$.
Consider a differentiable curve $\mathbf{c}(t)$ on the level set $L(k)$, that is $\mathbf{c}:I\to L(k)$ for some open interval $I\subseteq\mathbb{R}$. Consequently, we have that $g(\mathbf{c}(t))=k$ for all $t\in I$. So, by the chain rule, we obtain
$\left\langle\mathbf{c}^{\prime}(t),\nabla_{\mathbf{x}}g(\mathbf{c}(t))\right\rangle=0,$ (64)
[…]
We recall that the gradient is the direction of the greatest rate of increase of a function. Then, the source-like structures we observe on the plot of the vector field $P(\boldsymbol{\lambda})=-\nabla g(\boldsymbol{\lambda})$ (eg. see Figure 9) are the local maxima of the fidelity susceptibility $g$ in the direction transversal to the phase transition. The latter is consistent with the manifestation of QPT as a maxima of the susceptibility in finite-size systems.
Appendix I Scale invariance of the principal argument
We prove the scale-invariant property of the function $p(\boldsymbol{\lambda})$ defined in (13), which is employed in the construction of the phase diagrams.
We begin with a few concepts on the class of homogeneous functions. We call a function $f:K\subseteq\mathbb{R}^{n}\to\mathbb{R}$ positive homogeneous (PH) of degree $\alpha$ when
$f(t\mathbf{x})=t^{\alpha}f(\mathbf{x})\propto f(\mathbf{x})$ (65)
for all scaling factors $t>0$ such that $tx\in K$. The term ”positive” refers to the restriction on the scaling factor $t>0$, so the degree $\alpha$ can be any real number. The power law $x\mapsto c|x|^{\alpha}$ is PH. The parameter $\alpha$ is known as critical exponent in physics. Homogeneity is preserved by differentiation, so for $f:\mathbb{R}\to\mathbb{R}$
$\frac{\partial f(tx)}{\partial x}=$ $tf^{\prime}(tx)=t^{\alpha}f^{\prime}(x),$ (66)
(the right-most equality follows from the definition in (65)) dividing by $t$ we get $f^{\prime}(tx)=t^{\alpha-1}f^{\prime}(x)$ so $f^{\prime}$ is PH with degree $\alpha-1$. The latter shows that the vector field in (12) preserves PH. Instead, the angle of the vector field defined in (13) is scale-invariant (when the susceptibility follows the power law as we approach QPT). Indeed, by denoting the partial derivative $g_{\lambda_{i}}=\partial g/\partial\lambda_{i}$ of the function $g$ defined in (11), we have that
$p((\boldsymbol{\lambda}-\boldsymbol{\lambda}_{c})t)=$ $\arctan\left(\frac{t^{\alpha-1}g_{x_{2}}(\boldsymbol{\lambda}-\boldsymbol{\lambda}_{c})}{t^{\alpha-1}g_{x_{1}}(\boldsymbol{\lambda}-\boldsymbol{\lambda}_{c})}\right)=p(\boldsymbol{\lambda}-\boldsymbol{\lambda}_{c}),$ (67)
[…]
Appendix J Relation to information geometry
In this appendix, we discuss the fidelity susceptibilities relationship to information geometry and the Quantum Fisher Information (QFI). In (2) we introduced the Bures distance as the statistical distance between density matrices. The latter, when considered in terms of a small displacement $\mathrm{d}\boldsymbol{\lambda}$ in the vector of parameters $\boldsymbol{\lambda}$, induces the Bures metric [79]
$d^{2}_{B}(\rho(\boldsymbol{\lambda}),\rho(\boldsymbol{\lambda}+\mathrm{d}\boldsymbol{\lambda}))=$ $\sum_{i,j}h^{i,j}_{\boldsymbol{\lambda}}\mathrm{d}\lambda_{i}\,\mathrm{d}\lambda_{j},$ (68)
where $h_{\boldsymbol{\lambda}}$ is the metric tensor at $\boldsymbol{\lambda}$. In other words, the Bures metric is a notion of distinguishability of two density matrices that are infinitesimally close in the coordinate system $\boldsymbol{\lambda}$. In the realm of condensed matter physics, the Bures metric is better known as the fidelity susceptibility (introduced in Section II). Thus, the singularities of the metric $h_{\boldsymbol{\lambda}}$, correspond to the critical regions of the underlying model.
In quantum metrology, the quantum Cramér-Rao bound [80, 81], is the quantum equivalent to the classical lower-bound [82] of the variance of the unbiased estimator $\widehat{\lambda}$ (i.e. $\mathbb{E}[\widehat{\lambda}]=\lambda$) for an unknown parameter $\lambda$. The bound, in the case of a single parameter $\lambda$, is given by
$\mathrm{Var}(\widehat{\lambda})\geq\frac{1}{NQ(\rho(\lambda))},$ (69)
where $N$ is the number of independent repetitions and $Q(\rho(\lambda))$ is the QFI [79]. We consider a multi-parameter estimator $\widehat{\boldsymbol{\lambda}}$ the bound of the covariance matrix becomes
$\mathrm{Cov}(\widehat{\boldsymbol{\lambda}})\succcurlyeq\frac{Q(\rho(\boldsymbol{\lambda}))^{-1}}{N},$ (70)
[…]
There was consensus that, in the general case, the fidelity susceptibility is proportional to the quantum Fisher information [83], that is
$\mathcal{X}_{F}(\lambda)\overset{?}{=}$ $\frac{1}{4}Q(\rho(\lambda)).$ (71)
However, in [84] it is proved that at the points where we have a rank change of the density matrix $\rho(\lambda)$, the two quantities are not the same and the QFI presents a discontinuity. We note that the meaning and the consequent violation of the Cramér-Rao theorem on such singularities is still an active area of research [85, 86].
Concerning our formulation, the function $g(\boldsymbol{\lambda})$ defined in (11), can be related directly to the trace of the metric tensor, that is $g(\boldsymbol{\lambda})=-\operatorname{Tr}(h_{\boldsymbol{\lambda}})$. The minus factor is used as a convention to make the phase transitions appear as sources in the vector field resulting from (12). In regions of the parameter space $\mathcal{X}$ where the rank of $\rho(\boldsymbol{\lambda})$ is constant, it holds that $g(\boldsymbol{\lambda})$ is proportional to the trace of the QFI matrix. We also note that the trace of $h_{\boldsymbol{\lambda}}$ corresponds to the sum of the susceptibilities for each coupling parameter $\lambda_{i}$. The choice of the trace is also justified by the fact that this operator is basis-independent.
For a comprehensive introduction to information geometry and QPT, we refer to [87].